# Extensive entanglement between coupled Tomonaga-Luttinger liquids in and out of equilibrium

### v1

@@ 5 Summary and Outlook
Our study explores the dependence of entanglement and mutual information on key physical parameters, including temperature, tunneling strength, mean density, and the squeezing parameter. Our results suggest that mutual information—a key correlation measure employed in previous experiments [19, 25]—often reflects classical correlations and therefore cannot reliably capture the entanglement features of the system in the scenarios considered here. Nevertheless, it provides us with other important insights, including the quantum-to-classical correlation transition near the threshold temperature, and the conservation of mutual information in the non-interacting TLL dynamics, which forbids prethermalization to GGE states with distinct temperatures in the symmetric and antisymmetric sectors, unless phonon-phonon interaction is taken into account.
@@
A strong motivation for this work is the goal to experimentally detect extensive entanglement in many-body quantum systems. Our results suggest that such detection is not only feasible in 1D Bose gas experiments, but the required measurement protocol is also accessible. Specifically, one may apply the existing Gaussian tomography [23, 24] in both antisymmetric and symmetric sectors to reconstruct the full covariance matrix. This is made possible by recent developments in matter-wave interferometry after time-of-flight, allowing for single-shot extraction of not only relative but also common phase fluctuations [38]. Thus, while many studies on entanglement in field theories remain largely theoretical, this work highlights a clear pathway that could be pursued experimentally.
@@
On the other hand, experimental detection of entanglement brings about other challenges that have not been addressed here. For instance, certifying the reliability of Gaussian tomography reconstruction for witnessing entanglement includes taking into account statistical fluctuations, inherent systematic error in the phases extraction protocols [54, 38], and imperfections in the imaging which are known to introduce convolutions to the measured density interference patterns [52, 23]. Furthermore, typical experiments involve harmonic trapping, meaning that the system deviates from being fully homogeneous. In such situations, correlations between different momentum modes could become non-negligible, implying that logarithmic negativity is no longer a necessary and sufficient entanglement witness, and that the zero modes and the modes above the UV cutoff might affect the overall correlation. Even in experiments with a box potential [21, 18, 12, 19], the mean density curvature due to imperfect box should be scrutinized. A numerical and experimental study on precise certification of entanglement and mutual information from density interference data is an important direction for future work.
@@
Another direction of interest to both theory and experiment is to consider relaxation dynamics of entanglement and mutual information after a sudden quench (e.g. coherent splitting) or in the presence of a generic time-dependent tunneling strength. One may theoretically study these problems in the limit of low-energy theories—the Tomonaga-Luttinger liquid and the sine-Gordon model—or probe deviations from low energy theories and investigate the impact of cutoff and non-Gaussian correlations as discussed in the previous section. The latter is more challenging theoretically but it is highly relevant for experiments. Equipped with reliable Gaussian tomography, the long-time relaxation of Gaussian entanglement and mutual information can in principle be probed experimentally [77].
@@ 6 Acknowledgements
We are grateful to Jörg Schmiedmayer and Léonce Dupays for their valuable feedback. We also thank Paola Ruggiero, Julia Mathé, Nicky Kai Hong Li, and Giuseppe Vitagliano for insightful discussions. This work was supported by the start-up grant of the Nanyang Assistant Professorship of Nanyang Technological University, Singapore.
@@ Appendix A Derivation of logarithmic negativity and mutual information for tunnel-coupled gases in equilibrium
$=\frac{1}{2}\sqrt{(1+2\eta_{k}^{+})(1+2\eta_{k}^{-})}\left[\frac{\varepsilon_{k}^{-}}{\varepsilon_{k}^{+}}\frac{E_{k}}{E_{k}+2\hbar J}\right]^{\pm 1/2}.$ (37)
Entanglement is detected by logarithmic negativity if $\Gamma^{\top_{B}}$ violates the Heisenberg uncertainty principle, i.e. either $\nu_{k}^{+}<1/2$ or $\nu_{k}^{-}<1/2$ [See Eq. (16)]. Meanwhile, one can check from Eq. (37) that $\nu_{k}^{-}\geq 1/2$ by noticing that
$\left[\frac{\varepsilon_{k}^{-}}{\varepsilon_{k}^{+}}\frac{E_{k}}{E_{k}+2\hbar J}\right]^{-1/2}=\sqrt{\frac{(E_{k}+2\hbar J)(E_{k}+2gn_{\rm 1D})}{E_{k}(E_{k}+2\hbar J+2gn_{\rm 1D})}}\geq 1,$ (38)
@@
$=\frac{1}{4}\sqrt{\left[(1+2\eta_{k}^{+})+C_{k}(1+2\eta_{k}^{-})\right]\left[(1+2\eta_{k}^{+})+C_{k}^{-1}(1+2\eta_{k}^{-})\right]},$ (40)
where $C_{k}=(\varepsilon_{k}^{-}/\varepsilon_{k}^{+})(E_{k}/(E_{k}+2\hbar J))$. Then, for calculating the symplectic spectrum of the joint covariance matrix $\Gamma$, we perform the symplectic transformation (35) to $\Gamma$ and obtain the symplectic eigenvalues $\sigma_{k}^{\pm}$
@@ Appendix B Derivation of logarithmic negativity and mutual information for coherently split gases
Immediately after the coherent splitting, we assume thermal occupation in the symmetric sector and squeezed quantum noise in the antisymmetric sector [Eq. (26)]. The covariance matrix elements are now expressed as
@@
Meanwhile, the symplectic eigenvalues of $\Gamma^{aa}$ ($a\in\{A,B\}$) and $\Gamma$ are
$\lambda_{k}\Big{|}_{\Gamma^{aa}}=\frac{1}{4}\sqrt{\left(1+2\eta_{k}+\mathcal{C}_{k}\right)\left(1+2\eta_{k}+\mathcal{C}_{k}^{-1}\right)}\qquad\mathcal{C}_{k}=\frac{r^{2}}{2}\frac{\varepsilon_{k}}{E_{k}}$ (45)
$\sigma_{k}^{+}\Big{|}_{\Gamma}=\frac{1}{2}(1+2\eta_{k})\qquad\sigma_{k}^{-}\Big{|}_{\Gamma}=0$ (46)
From Eqs. (44)-(46) one can easily obtain the expression for logarithmic negativity and mutual information after coherent splitting [Eqs. (27) and (29)].
@@
Recall that entanglement occurs when the partially transposed covariance matrix violates the Heisenberg uncertainty priniciple, i.e., either $\nu_{k}^{+}<1/2$ or $\nu_{k}^{-}<1/2$. For the threshold temperature, we want to find a solution where at least one of them is exactly $1/2$.

### v2

@@ 5 Summary and Outlook
Our study explores the dependence of entanglement and mutual information on key physical parameters, including temperature, tunneling strength, mean density, and the squeezing parameter. Our analysis indicates that mutual information—a key correlation measure employed in previous experiments [19, 25]—often reflects classical correlations and therefore cannot reliably capture the entanglement features of the system in the scenarios considered here. Nevertheless, it provides us with other important insights, including the quantum-to-classical correlation transition near the threshold temperature, and the conservation of mutual information in the non-interacting TLL dynamics, which forbids prethermalization to GGE states with distinct temperatures in the symmetric and antisymmetric sectors, unless phonon-phonon interaction is taken into account.
@@
A strong motivation for this work is the goal to experimentally detect extensive entanglement in many-body quantum systems at finite temperatures. Our results suggest that such detection is not only feasible in 1D Bose gas experiments, but the required measurement protocol is also accessible. Specifically, one may apply the existing Gaussian tomography [23, 24] in both antisymmetric and symmetric sectors to reconstruct the full covariance matrix. This is made possible by recent developments in matter-wave interferometry after time-of-flight, allowing for single-shot extraction of not only relative but also common phase fluctuations [47]. Thus, while many studies on entanglement in field theories remain largely theoretical, this work highlights a clear pathway that could be pursued experimentally.
@@
On the other hand, experimental detection of entanglement brings about other challenges that have not been addressed here. For instance, certifying the reliability of Gaussian tomography reconstruction for witnessing entanglement includes taking into account statistical fluctuations, inherent systematic error in the phases extraction protocols [46, 47], and imperfections in the imaging, which are known to introduce convolutions to the measured density interference patterns [63, 23]. Furthermore, typical experiments involve harmonic trapping, meaning that the system deviates from being fully homogeneous. In such situations, correlations between different momentum modes can become non-negligible, so logarithmic negativity is no longer a necessary witness of entanglement. Additionally, contribution from the zero mode as well as from the modes above the UV cutoff may influence the overall correlations. Even in experiments with a box potential [21, 18, 12, 19], the mean density curvature due to imperfect box should be scrutinized. A numerical and experimental study on precise certification of entanglement and mutual information from density interference data is an important direction for future work.
@@
Another direction of interest to both theory and experiment is to consider relaxation dynamics of entanglement and mutual information after a sudden quench or in the presence of a generic time-dependent tunneling strength. One may theoretically study these problems in the limit of low-energy theories—the Tomonaga-Luttinger liquid and the sine-Gordon model—or probe deviations from low energy theories and investigate the impact of cutoff and non-Gaussian correlations as discussed in the previous section. The latter is more challenging theoretically but it is highly relevant for experiments. Equipped with reliable Gaussian tomography, the long-time relaxation of Gaussian entanglement and mutual information can in principle be probed experimentally [85].
@@ 6 Acknowledgements
We are grateful to Jörg Schmiedmayer and Léonce Dupays for their valuable feedback. We also thank Paola Ruggiero, Maximilian Prüfer, Giuseppe Vitagliano, Julia Mathé, Nicky Kai Hong Li, and Tiang Bi Hong for insightful discussions. This work was supported by the start-up grant of the Nanyang Assistant Professorship of Nanyang Technological University, Singapore.
@@ Appendix A Derivation of logarithmic negativity and mutual information for tunnel-coupled gases in equilibrium
$=\frac{1}{2}\sqrt{C_{k,J}^{\pm 1}(1+2\eta_{k}^{+})(1+2\eta_{k}^{-})},$ (38)
where $C_{k,J}$ is the spectral ratio defined in Eq. (22). Entanglement is detected by logarithmic negativity if $\Gamma^{\top_{B}}$ violates the Heisenberg uncertainty principle, i.e. either $\nu_{k}^{+}<1/2$ or $\nu_{k}^{-}<1/2$ [See Eq. (16)]. Meanwhile, one can check from Eq. (38) that $\nu_{k}^{-}\geq 1/2$ by noticing that
$C_{k,J}^{-1}=\left[\frac{\varepsilon_{k}^{-}}{\varepsilon_{k}^{+}}\frac{E_{k}}{E_{k}+2\hbar J}\right]^{-1}=\frac{(E_{k}+2\hbar J)(E_{k}+2gn_{\rm 1D})}{E_{k}(E_{k}+2\hbar J+2gn_{\rm 1D})}\geq 1,$ (39)
@@
$=\frac{1}{4}\sqrt{\left[(1+2\eta_{k}^{+})+C_{k,J}(1+2\eta_{k}^{-})\right]\left[(1+2\eta_{k}^{+})+C_{k,J}^{-1}(1+2\eta_{k}^{-})\right]},$ (41)
where $C_{k,J}=(\varepsilon_{k}^{-}/\varepsilon_{k}^{+})(E_{k}/(E_{k}+2\hbar J))$. Then, for calculating the symplectic spectrum of the joint covariance matrix $\Gamma$, we perform the symplectic transformation (36) to $\Gamma$ and obtain the symplectic eigenvalues $\sigma_{k}^{\pm}$
@@ Appendix B Derivation of logarithmic negativity and mutual information for coherently split gases
Immediately after the coherent splitting, we assume thermal occupation in the symmetric sector and quantum noise in the antisymmetric sector [Eq. (27)]. The covariance matrix elements are now expressed as
@@
Meanwhile, the symplectic eigenvalues of $\Gamma^{aa}$ ($a\in\{A,B\}$) are
$\lambda_{k}\Big|_{\Gamma^{aa}}=\frac{1}{4}\sqrt{\left(1+2\eta_{k}+\mathcal{C}_{k,r}\right)\left(1+2\eta_{k}+\mathcal{C}_{k,r}^{-1}\right)}\qquad\text{with}\qquad\mathcal{C}_{k,r}\coloneqq\frac{r^{2}}{2}\frac{\varepsilon_{k}^{+}}{E_{k}},$ (46)
and the symplectic eigenvalues of the joint covariance matrix $\Gamma$ are
$\sigma_{k}^{+}\Big|_{\Gamma}=\frac{1}{2}(1+2\eta_{k})\qquad\qquad\sigma_{k}^{-}\Big|_{\Gamma}=0.$ (47)
From the symplectic eigenvalues in Eqs. (45)-(47) one obtains the expression for logarithmic negativity and mutual information after coherent splitting [Eq. (28) and Eq. (30)].
@@
Recall that entanglement occurs when the partially transposed covariance matrix violates the Heisenberg uncertainty priniciple, i.e., either $\nu_{k}^{+}<1/2$ or $\nu_{k}^{-}<1/2$. In solving for the threshold temperature, we want to find a solution where at least one of them is exactly $1/2$.